On July 1, 2018, the Darden Company purchased a machine that was invoiced at $20,000 (terms 2/10, n/30, FOB destination). The invoice was paid within the discount period. The freight company charged the seller $500 to ship the machine, and the Darden Company incurred $800 in installation costs. The estimated useful life is five years, after which it will have an estimated salvage value of $1,000. Straight-line depreciation for 2018 would be:

A) $26,300 x 1/5 x 1/2 = $2,630.

B) $21,300 x 1/5 = $4,260.

C) $19,800 x 1/5 x 1/2 = $1,980.

D) $19,400 x 1/5 x 1/2 = $1,940.
Answer #1

Cost of the machine

= (20,000 - 2% discount) + installation 800

= 20,400

Depreciation = (Cost - Salvage value) /Useful life

= (20,400-1,000) * 1/5 * 6/12

= 19,400 * 1/5 * 6/12

= ​​​​​ 1940

Option D is the answer